OpenStudy (anonymous):

1/3x+1=1/2y ; Solve for y.

OpenStudy (anonymous):

2/3x+2=y

OpenStudy (turingtest):

answer or process?

OpenStudy (anonymous):

both i guess

OpenStudy (turingtest):

wait is it 1/3x or (1/3)x ?

OpenStudy (anonymous):

the second one.

OpenStudy (turingtest):

just multiply both sides by two: 2[(1/3)x+1]=2(1/2)y (2/3)x+2=y please be careful and use parentheses or you will get wrong answers from us